# Optimal Task Assignment and Path Planning using Conflict-Based Search with Precedence and Temporal Constraints

### 13 Feb 2024 v1

@@ 2. Background
In this section, we provide a brief overview of MAPF, CBS, MARL, and TAPF.
@@ 5.1. High-Level of CBS-TA-PTC
15:     P $\leftarrow$ node from OPEN with the highest return
@@
29:        Update Q solution for agent $a_{i}$ in constraint
@@ 5.2. Low-Level of CBS-TA-PTC
Multi-Label A* (MLA*) Grenouilleau et al. (2019) is used to generate a minimum sum of path costs solution for a task assignment under the constraints on the CT node. To speed up the low level, a linear programming module using one of the HiGHS solvers Hall et al. ([n.d.]); Huangfu and Hall (2018) is implemented to prune CT nodes that are unsolvable given their constraints from the various types of conflicts before the MLA* search as shown in lines 2426 in Algorithm 2.
@@ 6.1. Environment
Figure 1. An instance of the bomb-defusing task environment
@@ 6.2.2. Algorithm Setup
CBS-TA-PTC parameters are as follows: 1) Number of bombs per subtask and 2) Task allocation heuristic. The number of bombs per subtask is a proxy parameter for the number of goals per subtask in Algorithm 1, as the bomb sequence can be broken down into goals with defusing actions of a single color sequence. To keep the combinatorics of task allocation feasible, the number of bombs per subtask used will be in the range $[1,3]$. The task allocation heuristic is the heuristic used for HeuristicSort() at line 2 in Algorithm 1. We use the fuse length heuristic that sorts the bombs in ascending order of fuse length. For each combination of parameters, $5$ trials are executed with a timeout of $5$ minutes, and the success rate, optimality ratio, and runtime are recorded. Success is defined as finding an optimal solution before the timeout. The optimality ratio is the ratio between return and optimal return. The runtime includes only optimal solutions and all solutions (with cases due to timeout or no optimal solution found).
@@ 6.3. Results
6.3.1. MAPPO
Figure 2. Evaluated mean return of MAPPO from 16 rollouts
6.3.2. CBS
@@
Figure 3. Evaluated results for CBS-based algorithms

### 11 Apr 2024 v2

@@ 2. Background
In this section, we provide a brief overview of MAPF, CBS, MARL, and TAPF and highlight various related works to motivate and provide the necessary background for TAPF-PTC and CBS-TA-PTC.
@@ 4.1. High-Level of CBS-TA-PTC
15:     P $\leftarrow$ pop node from OPEN with the highest return
@@
29:        LowLevel() updates Q.solution for agent $a_{i}$ in constraint
@@ 4.2. Low-Level of CBS-TA-PTC
Multi-Label A* (MLA*) Grenouilleau et al. (2019) is used to generate a minimum sum of path costs solution for a task assignment under the constraints on the CT node as shown in line 5 and 29. To speed up the low level, a linear programming module using one of the HiGHS solvers Hall et al. ([n.d.]); Huangfu and Hall (2018) is implemented to prune CT nodes that are unsolvable given their constraints from the various types of conflicts before the MLA* search as shown in lines 2426 in Algorithm 2.
@@ 4.3. Theoretical Properties of CBS-TA-PTC
Figure 1. An instance of the bomb-defusing task environment
@@ 5.2.2. Algorithm Setup
CBS-TA-PTC parameters are as follows: 1) Number of bombs per subtask and 2) Task allocation heuristic. The number of bombs per subtask is a proxy parameter for the number of goals per subtask, $\beta$, in Algorithm 1, as the bomb sequence can be broken down into goals with defusing actions of a single color sequence. To keep the combinatorics of task allocation feasible, the number of bombs per subtask used will be in the range $[1,3]$. The task allocation heuristic is the heuristic used for HeuristicSort() at line 2 in Algorithm 1. We use the fuse length heuristic that sorts the bombs in ascending order of fuse length. For each combination of parameters, $10$ trials are executed with a timeout of $5$ minutes, and the success rate, optimality ratio, and runtime are recorded. Success is defined as finding an optimal solution before the timeout. The optimality ratio is the ratio between return and optimal return. The runtime includes only optimal solutions and all solutions (with cases due to timeout or no optimal solution found).
@@ 5.2.3. Hardware Setup
Figure 2. Evaluated mean return of MAPPO from 16 rollouts
@@ 5.3. Results
Figure 3. Evaluated results for CBS-based algorithms, where the intervals highlight the standard deviation of averaged results across 10 trials.
5.3.1. MAPPO
5.3.2. CBS